15. According to a recent press release, the UPS fleet includes more than 2,000 alternative fuel vehicles, in addition to its traditional fossil fuel truck lines. These alternative fuel vehicles deliver to homes and businesses every day alongside their traditional fuel counterparts. The alternative fleet includes compressed natural gas, propane-powered, hybrid electric, electric, liquefied natural gas, hydraulic hybrid, and fuel cell vehicles. This fictional problem outlines possible considerations. UPS has decided to replace a diesel-powered truck line in its fleet with electric hybrid drive train Class 5 trucks for environmental reasons (not cost savings). UPS has a choice between an Electric Vehicles International (EVI) electric hybrid and a similar truck manufactured by Efficient Drivetrains, Inc. (EDI). The EVI truck costs $130,000; saves $3,000 in fuel costs per year compared to the diesel truck; has a useful life of 4 years; and can be salvaged at a price of $30,000 (based on estimates for the price of a 4-year-old electric hybrid truck). The EDI truck costs $140,000; saves $4,000 in fuel costs per year; has a useful life of 5 years; and can then be salvaged at a price of $20,000 (based on estimates for the price of a 5-year-old electric hybrid truck). UPS depreciates its trucks according to 5-year MACRS, uses a WACC of 15% for ground transportation projects, and has a tax rate of 35%, working capital is unaffected by truck replacement. Assume inflation is 0% Which electric hybrid truck should UPS optimally select? Base your decision on an EAC analysis and assume the EAC for EDI is 27,093 a. EAC(EVI) 27,003 < EAC(EDI) 27,093, EVI is preferred b. EAC(EVI) 25,853 EAC(EDI) 27,093; EVI is preferred с. EAC(EVI-30,961 > EAC(EDI)-27,093; EDI is preferred d. EAC(EVI)-28,415 EAC(EDI) 27,093, EDI is preferred
Answer #1
EVI Trucks: 0 1 2 3 4 5
Initial cost -130000
After tax savings on fuel costs (3000*65%) 1950 1950 1950 1950
Depreciation 26000 41600 24960 14976 14976 7488 130000
Tax shield on depreciation at 35% 9100 14560 8736 5242
After tax salvage value [30000-(30000-22464)*0.35] 27362 22464
Net after tax annual cash flows -130000 11050 16510 10686 34554
PVIF at 15% 1 0.86957 0.75614 0.65752 0.57175 2.85498
PV at 15% -130000 9609 12484 7026 19756
NPV -81125
EAC = 81125/2.85498 = -28415
EDI Trucks: 0 1 2 3 4 5
Initial cost -140000
After tax savings on fuel costs (4000*65%) 2600 2600 2600 2600 2600
Depreciation 28000 44800 26880 16128 16128 8064 140000
Tax shield on depreciation at 35% 9800 15680 9408 5645 5645
After tax salvage value [20000-(20000-8064)*0.35] 15822
Net after tax annual cash flows -140000 12400 18280 12008 8245 24067
PVIF at 15% 1 0.86957 0.75614 0.65752 0.57175 0.49718 3.35216
PV at 15% -140000 10783 13822 7895 4714 11966
NPV -90820
EAC =102786/3.35216 = -27093
ANSWER: OPTION [d] EAC(EVI) 28,415>EAC(EDI) = 27093: EDI IS PREFERRED.